OpenStudy (anonymous):

PLEASE HELP Write a sine function with a maximum of 0.5, a minimum of -0.5, and a period of 2(pi)/5. WILL FAN AND MEDAL~ OuO

OpenStudy (anonymous):

@dan815

OpenStudy (johnweldon1993):

Hey again :D So we know the basic equation for a sine function is \[\large Asin(Bx - C) + D\] A is the amplitude (how high up and down the wave reaches) B is the period (how often the wave repeats) C is the phase shift And D is the vertical displacement

OpenStudy (johnweldon1993):

So A...the amplitude will be the 0.5 that you want the sine wave to reach... If the amplitude is the 0.5....the sine wave will oscillate between 0.5 and -0.5

OpenStudy (johnweldon1993):

And the B is our period...now the period of a sine wave is NORMALLY \(\large 2\pi \) However the period can be calculated by using \[\large Period = \frac{2\pi}{|B|}\] Now here, since you're given the period to be \(\Large \frac{2\pi}{5}\) So \[\large Period = \frac{2\pi}{B} = \frac{2\pi}{5}\] So we can see our B = 5

OpenStudy (johnweldon1993):

Now C and D we need not worry about...in this case they are both 0... So altogether, our sine wave is \[\large 0.5sin(5x)\]

OpenStudy (anonymous):

sorry for the late responce!

OpenStudy (anonymous):

thanks for explaining the formula, i can actually follow that!